# Retrieval Enhanced Feedback via In-context Neural Error-book

### v2

@@ Abstract
Jongyeop Hyun School of CSE Chung-Ang University jesussuej@cau.ac.kr mldljyh@postech.ac.kr Bumsoo Kim School of CSE Chung-Ang University bumsoo@cau.ac.kr
@@ 3 Method
The objective of this paper is to develop a structured Neural Error-book that systematically provides feedback to enhance model performance effectively. To achieve this, the Neural Error-book construction involves three distinct feedback formulation stages.
@@ 3.2 Feedback Filtering
After Feed-Target, Feed-Check and Feed-Path, the resulting feedback instances are categorized as either Task/Process-relative feedback (that directly corrects task-specific errors or adds specifics for reasoning, e.g., “Adjust counts for occluded objects.”, “Verify spatial relationships again.”) and Self-Regulatory feedback (that addresses metacognitive habits or personal traits, e.g., “To improve accuracy, try solving similar problems multiple times.”, “You are good at object recognition.”). Empirical observations suggest that self-regulatory feedback tends to hinder final Chain-of-Thought (CoT) performance. Therefore, we classify the resulting feedback instances and filter out self-regulatory feedback that does not contribute positively.

### v3

@@ Abstract
Jongyeop Hyun School of CSE Chung-Ang University jesussuej@cau.ac.kr††thanks: contact: mldljyh@postech.ac.kr Bumsoo Kim School of CSE Chung-Ang University bumsoo@cau.ac.kr
@@ 3 Method
The objective of this paper is to develop a structured Neural Error-book that systematically provides feedback to enhance model performance effectively. To achieve this, the Neural Error-book construction involves three distinct feedback formulation stages. The detailed pseudocode for our REFINE framework is presented in Appendix B.
@@ 3.2 Feedback Filtering
After Feed-Target, Feed-Check and Feed-Path, the resulting feedback instances are categorized. This classification is performed automatically by our teacher model. The model is provided with the definitions of “Task/Process-relative” feedback (that directly corrects task-specific errors or adds specifics for reasoning, e.g., “Adjust counts for occluded objects.”) and “Self-Regulatory” feedback (that addresses metacognitive habits or personal traits, e.g., “To improve accuracy, try solving similar problems multiple times.”) from Hattie and Timperley (2007) and then prompted to classify each generated instance. This automated method ensures the process is consistent and reproducible. Based on empirical observations suggesting that self-regulatory feedback tends to hinder final Chain-of-Thought (CoT) performance, we filter out feedback classified as self-regulatory.
@@ 5 Conclusion
Appendix A Dataset Statistics
| Benchmark | Error-book Size | Test Size |
| MME-RealWorld (Reasoning) | 534 | 2,092 |
| MMStar (Reasoning) | 123 | 250 |
| SEED-Bench-2-Plus | 563 | 1,141 |
The Neural Error-book for each benchmark was constructed using only the questions the student model answered incorrectly during the training phase.
For the MME-RealWorld benchmark, the Error-book was created based on errors from the MME-RealWorld-Lite subset.
For MMStar and SEED-Bench-2-Plus, the datasets were split into equal halves to create distinct train and test sets. The Error-book for these benchmarks was then built using the incorrectly answered questions from their respective 50% train sets.